if 3,9,27 is a geometric sequence with common ratio of 3. (a)find the value of 10term (b)sum of first 15term
(2)suppose we have 8,16,32,64. (a)find the value of the 20term (b)sum of the first 15term

Answer by stanbon(75887)

if 3,9,27 is a geometric sequence with common ratio of 3.
a(n) = a(1)*r^(n-1)
-------------------------------
(a)find the value of 10 term
a(10) = 3*3^9 = 3^10
----------------------------
(b)sum of first 15 term
S(n) = a[(1-r^n)/(1-r)]
S(15) = 3[(1-3^15)/(-2)] = 21,523,359
------------------------------------
(2)suppose we have 8,16,32,64.
(a)find the value of the 20term
(b)sum of the first 15term
===
Use the formulas.
=========================
Cheers,
Stan H.
==========